# Tailoring Chatbots for Higher Education: Some Insights and Experiences

### v1

@@ Acknowledgements.
- OpenAI [2024a] OpenAI, GPT-4, https://openai.com/gpt-4 (2024a).
- Google [2024] Google, Gemini, https://gemini.google.com/ (2024), accessed: 2024-08-22.
- Meta [2024] Meta, Llama 3, https://llama.meta.com/docs/model-cards-and-prompt-formats (2024).
- Apple [2024] Apple, Apple Intelligence, https://www.apple.com/apple-intelligence/ (2024), accessed: 2024-08-22.
- Mistral AI [2024] Mistral AI, Mistral, https://mistral.ai (2024), accessed: 2024-08-22.
- Anthropic [2024] Anthropic, Claude, https://claude.ai/ (2024), accessed: 2024-08-22.
- OpenAI [2024b] OpenAI, Hello GPT-4o, https://openai.com/index/hello-gpt-4o/ (2024b).
@@
- ETH AI Center & EPFL AI Center [2024] ETH AI Center & EPFL AI Center, SWISS AI Initiative, https://www.swiss-ai.org (2024), accessed: 2024-08-22.
@@
- Achiam et al. [2023] J. Achiam, S. Adler, S. Agarwal, L. Ahmad, I. Akkaya, F. L. Aleman, D. Almeida, J. Altenschmidt, S. Altman, S. Anadkat, et al., GPT-4 technical report, arXiv preprint arXiv:2303.08774 (2023).
- OpenAI [2024c] OpenAI, Fine-tuning, https://platform.openai.com/docs/guides/fine-tuning (2024c), accessed: 2024-08-22.
@@
- OpenAI [2024e] OpenAI, OpenAI Embeddings, https://platform.openai.com/docs/guides/embeddings (2024e), accessed: 2024-08-22.
- Kortemeyer [2024a] G. Kortemeyer, Embeddings and RAG with Azure OpenAI API, https://cscblog.ethz.ch/index.php/2024/02/06/az-open-ai-rag-chromadb-langchain/ (2024a), accessed: 2024-08-22.
- Kortemeyer [2024b] G. Kortemeyer, Ethel chat source code, https://gitlab.ethz.ch/ethel/etheltest/-/tree/main/rag (2024b), accessed: 2024-07-28.
- Kortemeyer [2024c] G. Kortemeyer, Ethel: A virtual teaching assistant, arXiv preprint arXiv:2407.19452 (2024c).
- OpenAI [2024f] OpenAI, Introducing ChatGPT Edu, https://openai.com/index/introducing-chatgpt-edu/ (2024f).
- Pham et al. [2024] D. Pham, J. Sheffey, C. M. Pham, and A. Houmansadr, Proxygpt: Enabling anonymous queries in AI chatbots with (un)trustworthy browser proxies, arXiv preprint arXiv:2407.08792 (2024).
- Hochschule für angewandte Wissenschaft und Kunst Hildesheim/Holzminden/Göttingen [2024] Hochschule für angewandte Wissenschaft und Kunst Hildesheim/Holzminden/Göttingen, HAWKI, https://www.swiss-ai.org (2024), accessed: 2024-08-22.
- Microsoft [2024] Microsoft, Azure AI services, https://azure.microsoft.com/en-us/products/ai-services (2024), accessed: 2024-08-22.
- Amazon Web Services [2024] Amazon Web Services, Amazon SageMaker Model Deployment, https://aws.amazon.com/sagemaker/deploy/ (2024), accessed: 2024-08-22.

### v2

@@ Acknowledgements.
- Muscanell et al. [2025] N. Muscanell, J. Petko, and Y. Chan, Ai and tech on campus: Real talk with students and faculty, EDUCAUSE Review (Online) (2025).
- Brodheim [2025] T. Brodheim, Shaping the future of learning: Ai in higher education, EDUCAUSE Review (Online) (2025).
- Radford et al. [2018] A. Radford, K. Narasimhan, T. Salimans, and I. Sutskever, Improving language understanding by generative pre-training, https://cdn.openai.com/research-covers/language-unsupervised/language_understanding_paper.pdf (2018), foundational report introducing the GPT approach.
- Gemini Team, Google [2023] Gemini Team, Google, Gemini: A family of highly capable multimodal models, arXiv 10.48550/arXiv.2312.11805 (2023), arXiv:2312.11805 [cs.CL] .
- Touvron et al. [2023] H. Touvron, T. Lavril, G. Izacard, X. Martinet, M.-A. Lachaux, T. Lacroix, B. Rozière, N. Goyal, E. Hambro, F. Azhar, A. Rodriguez, A. Joulin, E. Grave, and G. Lample, Llama: Open and efficient foundation language models, arXiv 10.48550/arXiv.2302.13971 (2023), arXiv:2302.13971 [cs.CL] .
- Swiss AI Initiative [2025a] Swiss AI Initiative, Apertus, https://www.swiss-ai.org/apertus (2025a), open multilingual foundation model; model cards and documentation.
- Swiss AI Initiative [2025b] Swiss AI Initiative, Swiss ai initiative — overview, https://www.swiss-ai.org/ (2025b), open science/open source effort for AI foundation models.
- TOP500 Project [2025] TOP500 Project, Alps — system #180259, https://top500.org/system/180259/ (2025), rank 8 on the TOP500 June 2025 list.
- OpenAI [2025a] OpenAI, GPT-5 System Card, Tech. Rep. (OpenAI, 2025) model overview and safety documentation.
- OpenAI [2025b] OpenAI, Introducing GPT-5, https://openai.com/index/introducing-gpt-5/ (2025b).
@@
- OpenAI [2024a] OpenAI, Hello GPT-4o, https://openai.com/index/hello-gpt-4o/ (2024a).
@@
- OpenAI [2024b] OpenAI, Fine-tuning, https://platform.openai.com/docs/guides/fine-tuning (2024b), accessed: 2024-08-22.
@@
- Pal Chowdhury et al. [2024] S. Pal Chowdhury, V. Zouhar, and M. Sachan, Autotutor meets large language models: A language model tutor with rich pedagogy and guardrails, in Proceedings of the Eleventh ACM Conference on Learning@ Scale (2024) pp. 5–15.
- OpenAI [2024c] OpenAI, OpenAI Embeddings, https://platform.openai.com/docs/guides/embeddings (2024c), accessed: 2024-08-22.
- Kortemeyer [2024a] G. Kortemeyer, Ethel chat source code, https://gitlab.ethz.ch/ethel/etheltest/-/tree/main/rag (2024a), accessed: 2024-07-28.
- Kortemeyer [2024b] G. Kortemeyer, Ethel: A virtual teaching assistant, arXiv preprint arXiv:2407.19452 (2024b).
- Geisler and Kortemeyer [2025] T. Geisler and G. Kortemeyer, When ai evaluates its own work: Validating learner-initiated, ai-generated physics practice problems, arXiv preprint arXiv:2508.03085 (2025).
- Kortemeyer and Nöhl [2025] G. Kortemeyer and J. Nöhl, Assessing confidence in ai-assisted grading of physics exams through psychometrics: An exploratory study, Physical Review Physics Education Research 21, 010136 (2025).
- Cvengros and Kortemeyer [2025] J. Cvengros and G. Kortemeyer, Assisting the grading of a handwritten general chemistry exam with artificial intelligence, arXiv preprint arXiv:2509.10591 (2025).
- Kortemeyer et al. [2025] G. Kortemeyer, A. Caspar, and D. Horica, Artificial-intelligence grading assistance for handwritten components of a calculus exam, arXiv preprint arXiv:2510.05162 (2025).